OpenStudy (anonymous):

A deck of cards contains 5 black cards numbered 5 to 9 and 5 red cards numbered 1 to 5. How many 6-card hands having 4 black cards and 2 red cards can be dealt? 15 50 25 300

OpenStudy (bibby):

You have 10 total cards, and you're getting a 6 card hand. Since you're trying to do 4 blacks and 2 reds, that's 75% black, 5^2 is 25

OpenStudy (anonymous):

Thank you for your help guys!

OpenStudy (perl):

the numerator is the number of ways (woops i did the probability of dealing that hand)

OpenStudy (anonymous):

Thats okay, thank you! Just double checking are the numbers inside the parenthesis being divided?

OpenStudy (perl):

the number of ways is 5 choose 4 * 5 choose 2 = 5 * 10 = 50

OpenStudy (perl):

$$ \Large \rm { {5 \choose 4 }\cdot {5 \choose 2 } }$$

OpenStudy (perl):

that is a special notation

OpenStudy (anonymous):

Ohhh okay gotcha, Thank you!

OpenStudy (perl):

also im assuming the order does not count how you deal the cards. what matters is that you have 4 black and 2 red

OpenStudy (anonymous):

That sounds about right. Thanks!